OpenStudy (amistre64):

Does \(\int \frac{1}{x+i}=ln|x+i|\), or is there something peculiar about using complex numbers?

OpenStudy (amistre64):

i spose i shoulda included a "dx" :)

OpenStudy (anonymous):

That should be right, any integral of 1/# will be natural log with or without complex numbers. As long as there isn't any multiplying or anything.

OpenStudy (anonymous):

i suppose no ... it is just treated as constant

OpenStudy (amistre64):

if we decompose a fraction into a complex linear; how would that affect the outcome of the integration?

OpenStudy (amistre64):

\[\int\frac{1}{x^2+1}dx=tan^{-1}(x)+C\] but what if we decompose it?

OpenStudy (zarkon):

you could also do this... \[\int\frac{1}{x+i}dx=\int\frac{1}{x+i}\frac{x-i}{x-i}dx\] \[=\int\frac{x-i}{x^2+1}dx=\ln(x^2+1)-\tan^{-1}(x)i+c\]

OpenStudy (zarkon):

oops \[=\int\frac{x-i}{x^2+1}dx=\frac{\ln(x^2+1)}{2}-\tan^{-1}(x)i+c\]

OpenStudy (anonymous):

@zarkon : its\[\frac{\ln(x ^{2}+1)}{2} - i (\tan^{-1} (x ))+c\]

OpenStudy (zarkon):

I know